# subequation

### v1

@@ 4 Conclusions
We have presented the leading order coupling in BOEFT of double-heavy hadrons to pairs of heavy hadrons for any spin of the light degrees of freedom. We have shown how to transform the expression of the BOEFT states from using a basis of fields in a standard spherical spin representation to a basis of fields belonging to $D_{\infty h}$ representations. In this base the coupling becomes extremely simple and the computation of the expected value between a double-heavy hadron and a heavy hadron pair states is straightforward. Using the expected value of the coupling, the contribution of the heavy hadron pairs to the self-energy of the double-heavy hadrons can be computed, and from it the effect of the former on the mass and width of the latter. The matching of the potential accompanying of the coupling operator to an NRQCD correlator has not been presented. However, its can be straightforwardly obtained following the matching computation for the coupling potential for conventional quarkonium to a heavy-meson pair presented in Ref. TarrusCastella:2022rxb , by replacing the quarkonium interpolating operator for the double-heavy hadron one, which can be found in Ref. Soto:2020xpm .
@@
In section 3 we have applied the results of section 2 for the case of the coupling of $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ to hybrid quarkonium. In particular we have studied the hybrid states generated by the two lowest static energies, $\Sigma_{u}^{-}$ and $\Pi_{u}$. From the computation of the expected value of the coupling operator we have found selection rules for these decays. The most important one is that a hybrid state without a $\Sigma_{u}^{-}$ component cannot decay through the leading order coupling operator. Secondly, only a single angular momentum, equal to the hybrid total angular momentum quantum number $\ell$, is allowed for the heavy meson pair product of the decay. Based on these selection rules and the leading order masses, we identify the $1p_{1}$ state, both the in charm and bottom sectors, as having very suppressed decays into any heavy-meson pairs. Two of the states in the corresponding heavy quark spin-symmetry multiplets have the explicitly exotic $j^{pc}$, thus being good candidates to be unambiguously identified as hybrids.
@@
The potential in the coupling operator must be computed with nonperturbative techniques and so far is unknown. We have built a model for it based on its expected short- and long distance behaviors and inspired by the analogous potential coupling conventional quarkonium to heavy meson pairs for which there exists lattice data. For three variations of this model potential we have obtained the energy shifts caused by the heavy meson pairs to the hybrid masses as well as the decay widths of the hybrid states into the heavy meson pairs. We presented our results in tables 4 and 5. A remarkable result is that the energy shifts and widths are larger in the bottomonium sector than in the charmonium one. This is a result of the coupling operator in Eq. (9) not being heavy-quark mass suppressed and the potentials (in Fig. 2) having the larger values in the short-distance. A comparison with the experimental quarkonium states above open flavor threshold can be found in Figs. 3 and 4. In the charmonium sector $\psi(4160)$, $\psi(4415)$, $\chi_{c1}(4274)$ and $\chi_{c0}(4700)$ remain good candidates for hybrid states. In the bottomonium sector the hybrid states turn out to have too large widths to be compatible experimental states observed so far.

### v2

@@ 4 Conclusions
We have presented the leading order BOEFT couplings of double-heavy hadrons to pairs of heavy hadrons for any spin of the light degrees of freedom. We have shown how to transform the expression of the BOEFT states from using a basis of fields in a standard spherical spin representation to a basis of fields belonging to $D_{\infty h}$ representations. In this base the coupling becomes extremely simple and the computation of its expected value between a double-heavy hadron and a heavy hadron pair states is straightforward. Using the expected value of the coupling, the contribution of the heavy-hadron pairs to the self-energy of the double-heavy hadrons can be computed, and from it the effect of the former on the mass and width of the latter. The matching of the potential accompanying the coupling operator to a NRQCD correlator has not been presented. However, it can be straightforwardly obtained following the matching computation for the coupling potential for conventional quarkonium to a heavy-meson pair presented in Ref. TarrusCastella:2022rxb , by replacing the quarkonium interpolating operator for the double-heavy hadron one, which can be found in Ref. Soto:2020xpm .
@@
In Section 3 we have applied the results of Section 2 for the case of the coupling of $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ to hybrid quarkonium. In particular, we have studied the hybrid states generated by the two lowest static energies, $\Sigma_{u}^{-}$ and $\Pi_{u}$. From the computation of the expected value of the coupling operator we have found selection rules for these decays. The first one is that a hybrid state without a $\Sigma_{u}^{-}$ component cannot decay through the leading order coupling operator. Secondly, only a single angular momentum, equal to the hybrid total angular momentum quantum number $\ell$, is allowed for the heavy-meson pair product of the decay. The $1p_{1}$ states, both the in charm and bottom sectors, turn out to be very interesting. Due to being pure $\Pi_{u}$, these states have $1/m_{Q}$-suppressed decays to $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ and, due to their leading order mass, they are not kinematically allowed to decay to any other heavy-meson pair. Furthermore, two of the $j^{pc}$ states in the corresponding heavy quark spin-symmetry multiplets are explicitly exotic, thus being good candidates to be unambiguously identified as hybrids.
@@
The potential in the coupling operator must be computed with nonperturbative techniques and so far is unknown. We have built a model for it based on its expected short- and long-distance behaviors and inspired by the analogous potential coupling conventional quarkonium to heavy-meson pairs for which there exists lattice data. For three variations of this model potential we have obtained the energy shifts caused by the heavy-meson pairs to the hybrid masses as well as the decay widths of the hybrid states into the heavy-meson pairs. We presented our results in tables 4 and 5. A remarkable result is that the energy shifts and widths are larger in the bottomonium sector than in the charmonium one. This is a result of the coupling operator in Eq. (9) not being heavy-quark mass suppressed and the potentials (in Fig. 2) having the larger values in the short-distances. A comparison with the experimental quarkonium states above open flavor threshold can be found in Figs. 3 and 4. In the charmonium sector $\psi(4160)$, $\psi(4415)$, $\chi_{c1}(4274)$ and $\chi_{c0}(4700)$ remain good candidates for hybrid states. In the bottomonium sector the hybrid states turn out to have too large widths to be compatible with the experimental states observed so far.